Fred gave one half of his baseball cards to Sally. Sally gave Jeff half of her cards that she got for Fred. Jeff gave Allen half of the cards that he got from Sally. Allen got 6 cards. How many cards did Fred originally have?
Answer by rfer(16322)

6/(1/8)=
6*8/1=48